A ceiling fan with 80-cm-diameter blades is turning at 60 rpm.Suppose the fan coasts to a stop 25s after being turned off.

a.) What is the speed of the trip of a blade 10 s after the fan is turned off?

b.) Through how many revolutions does the fan turn while stopping?
Answer #1
Concepts and reason

The required concepts to solve these questions are rotational kinematic equations, angular velocity, angular acceleration and angular displacement.

Firstly, calculate the angular acceleration by using angular velocity equation. Calculate the linear speed of the trip of a blade 10 s after the fan is turned off by using angular velocity equation. Calculate the number of revolutions by the fan while stopping by using angular displacement equation.

Fundamentals

The expression for angular velocity is expresses as follows,

ω=ω0+αt\omega = {\omega _0} + \alpha t

Here, α\alpha is the angular acceleration, tt is the time and ω0{\omega _0}is the initial angular velocity.

The expression for angular displacement as a function of time is expresses as follows,

θ=ω2ω022α\theta = \frac{{{\omega ^2} - {\omega _0}^2}}{{2\alpha }}

Here, ω0{\omega _0}is the initial angular velocity, ω\omega is final angular velocity. and α\alpha is the angular acceleration.

The expression for linear speed in term of angular velocity is expresses as follows,

v=rωv = r\omega

Here, rris the radius and ω\omega is the angular velocity.

The expression for radius is expresses as follows,

r=d2r = \frac{d}{2}

Here, dd is the diameter.

(a)

The expression for angular velocity is expresses as follows,

ω=ω0+αt\omega = {\omega _0} + \alpha t …… (1)

Rearrange the equation for angular acceleration.

α=ωω0t\alpha = \frac{{\omega - {\omega _0}}}{t} …… (2)

Convert the unit of (ω0)\left( {{\omega _0}} \right)initial angular velocity (60rpm)\left( {60{\rm{ rpm}}} \right)torads1{\rm{rad}} \cdot {{\rm{s}}^{ - 1}}.

60rpm=(60rpm)(0.104rads11rpm)=6.24rads1\begin{array}{c}\\60{\rm{ rpm}} = \left( {60{\rm{ rpm}}} \right)\left( {\frac{{0.104{\rm{ rad}} \cdot {{\rm{s}}^{ - 1}}}}{{1{\rm{ rpm}}}}} \right)\\\\ = 6.24{\rm{ rad}} \cdot {{\rm{s}}^{ - 1}}\\\end{array}

Substitute 00 for ω\omega , 6.24rads16.24{\rm{ rad}} \cdot {{\rm{s}}^{ - 1}} for ω0{\omega _0} and25s25{\rm{ s}} for ttin the equation (2).

α=0(6.24rads1)(25s)=0.24rads2\begin{array}{c}\\\alpha = \frac{{0 - \left( {6.24{\rm{ rad}} \cdot {{\rm{s}}^{ - 1}}} \right)}}{{\left( {25{\rm{ s}}} \right)}}\\\\ = - 0.24{\rm{ rad}} \cdot {{\rm{s}}^{ - 2}}\\\end{array}

Substitute 6.24rads16.24{\rm{ rad}} \cdot {{\rm{s}}^{ - 1}} for ω0{\omega _0} , 0.24rads2 - 0.24{\rm{ rad}} \cdot {{\rm{s}}^{ - 2}} for α\alpha and10s10{\rm{ s}} for ttin the equation (1).

ω=6.24rads1+(0.24rads2)(10s)=3.84rads1\begin{array}{c}\\\omega = 6.24{\rm{ rad}} \cdot {{\rm{s}}^{ - 1}} + \left( { - 0.24{\rm{ rad}} \cdot {{\rm{s}}^{ - 2}}} \right)\left( {10{\rm{ s}}} \right)\\\\ = 3.84{\rm{ rad}} \cdot {{\rm{s}}^{ - 1}}\\\end{array}

The expression for radius is,

r=d2r = \frac{d}{2}

Substitute 80cm80{\rm{ cm}} for ddin the above equation.

r=80cm2=40cm(1m102cm)=40×102m\begin{array}{c}\\r = \frac{{80{\rm{ cm}}}}{2}\\\\ = 40{\rm{ cm}}\left( {\frac{{1{\rm{ m}}}}{{{{10}^2}\,{\rm{cm}}}}} \right)\\\\ = 40 \times {10^{ - 2}}\;{\rm{m}}\\\end{array}

The expression for linear speed of the trip of a blade 10 s after the fan is turned off,

v=rωv = r\omega

Substitute 3.84rads13.84{\rm{ rad}} \cdot {{\rm{s}}^{ - 1}} for ω\omega and 40×102m40 \times {10^{ - 2}}\;{\rm{m}}for rrin the above equation.

v=(3.84rads1)(40×102m)=1.536ms1\begin{array}{c}\\v = \left( {3.84{\rm{ rad}} \cdot {{\rm{s}}^{ - 1}}} \right)\left( {40 \times {{10}^{ - 2}}\;{\rm{m}}} \right)\\\\ = 1.536{\rm{ m}} \cdot {{\rm{s}}^{ - 1}}\\\end{array}

(b)

The expression for angular displacement as a function of time is,

θ=ω2ω022α\theta = \frac{{{\omega ^2} - {\omega _0}^2}}{{2\alpha }}

Substitute 00 for ω\omega , 6.24rads16.24{\rm{ rad}} \cdot {{\rm{s}}^{ - 1}} for ω0{\omega _0} , 0.24rads2 - 0.24{\rm{ rad}} \cdot {{\rm{s}}^{ - 2}} for α\alpha in the above equation.

θ=0(6.24rads1)22(0.24rads2)=81.12rad(1rev2πrad)=12.91rev\begin{array}{c}\\\theta = \frac{{0 - {{\left( {6.24{\rm{ rad}} \cdot {{\rm{s}}^{ - 1}}} \right)}^2}}}{{2\left( { - 0.24{\rm{ rad}} \cdot {{\rm{s}}^{ - 2}}} \right)}}\\\\ = 81.12{\rm{ rad}}\left( {\frac{{\;1{\rm{ rev}}}}{{2\pi {\rm{ rad}}}}{\rm{ }}} \right)\\\\ = 12.91{\rm{ rev}}\\\end{array}

Ans: Part a

The speed of the trip of a blade 10 s after the fan is turned off is 1.536ms11.536{\rm{ m}} \cdot {{\rm{s}}^{ - 1}}.

Part b

The number of revolutions by the fan while stopping is 12.91rev12.91{\rm{ rev}}.